Blue Ridge Marketing Inc. manufactures two products, A and B. Presently, the company uses a single plantwide factory overhead rate for allocating overhead to products. However, management is considering moving to a multiple department rate system for allocating overhead. The following table presents information about estimated overhead and direct labor hours. Overhead Direct Labor Hours (dlh) Product A B Painting Dept. $232,900 8,200 dlh 13 dlh 3 dlh Finishing Dept. 71,100 9,200 2 18 Totals $304,000 17,400 dlh 15 dlh 21 dlh The factory overhead allocated per unit of Product B in the Painting Department if Blue Ridge Marketing Inc. uses the multiple production department factory overhead rate method is a. $23.18 per unit b. $85.20 per unit c. $28.40 per unit d. $52.41 per unit
Answer #1
Overhead allocation rate for painting department=estimated overhead/estimated direct labor hours=232900/8200=$ 28.40 per direct labor hour
Factory overhead allocated to product B=DL hours used by B in painting*overhead allocation rate=3*28.40=$ 85.20 per unit
Answer is
b. $ 85.20 per unit